A particle is thrown vertically upwards from the surface of the earth and it reaches to a maximum height equal to the radius of earth.The ratio of the velocity of projection to the escape velocity on the surface of earth is

Grade:11

1 Answers

Mohammed Kasib Siddiqui
5 years ago
W= mgh/(1+h/R) which is change in potential also .
Apply law of conservation of mass.
KE=PE
1/2mv^2= mgh/(1+h/R)
where h=R
On solving we get v= root (gR)
Escape velocity = root (2gR)
Now divide these velocities to get yoir answer
 
Plz approve.